# Entanglement teleportation along a regenerating hamster-wheel graph state

### v2

@@ Abstract
Figure 3: Polar plots of the teleported two-qubit graph state showing (a) negativity and (b) fidelity on the radial axis versus the number of teleportation hops in the angular axis sampled from the H1-1 emulator averaged over four trials. The results for larger than 75 hops are marked using different colours to highlight the glitch in the measured entanglement due to different noise parameters resulting from a long time gap between experiments. The three data points marked in red represent results obtained from the real H1-1 quantum computer for 9, 18 and 56 hops with respective negativities $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelities $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$, where experiments were conducted on 21 March, 25 March and 24 April 2024 respectively. Assuming the distribution is Gaussian, the half-width of the shaded strip is two times the standard error of the sampled negativity or fidelity to represent 95% confidence interval. On the other hand, the data obtained from real H1-1 quantum computers are only sampled in a single trial due to limited resources. Therefore, we use the technique of bootstrapping to resample the data from the measured sample distribution for 200 trials. The error bars represent 95% confidence interval of the bootstrapped negativities or fidelities. All results are mitigated with REM as indicated in Teleportation quality witnesses. Due to the feedforward gates applied based on mid-circuit measurements in the dynamic circuit approach, REM is applied to only the two qubits of the final teleported state. The corresponding plot with regular crossing axes are included in Figure 4.
@@
In Figure 3a and Figure 3b, we present the negativities and fidelities recorded from the real machine (red-marked data points) and the emulator (blue data), plotted in polar plots against the number of hops that the two-qubit graph states were teleported across using the teleportation strategies. The emulated data is averaged over four trials. Due to limited access to computational resources, experiments on the H1-1 quantum computer were conducted for only three scenarios of 9, 18, and 56 hops in a single trial of the QST experiment with 1000 shots for each QST circuit. We bootstrap 200 samples from each QST basis measurement for both negativities and fidelities. The error bars of the data obtained from the H1-1 quantum computer of size $\epsilon$ take the half-range of the 95% confidence interval of the bootstrapped negativities or fidelities, such that $\epsilon=\frac{1}{2}{\left(\left\lfloor a_{0.975N}\right\rfloor-\left\lceil a_{0.025N}\right\rceil\right)}$, where $\{a_{n}\},n=1,2,\cdots,N$ is the sequence of bootstrapped negativities or fidelities ordered from lowest to highest and $N=200$ is the sample size. For both the negativity and fidelity, we have used REM, the nearest physical probability vector, and the nearest physical density matrix algorithm as previously outlined to mitigate classical readout errors. After all, from the H1-1 quantum computer we obtain the negativity of $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelity of $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$ corresponds to 9, 18 and 56 hops respectively. For the data obtained from the emulator, we report a clear trend that the entanglement of the teleported state decreases gradually, almost linearly, as the number of hops increases. This trend is closely in agreement with the data points obtained from the real device without significant deviations, as well as the trend of the two-qubit gate fidelity plot in the H1 model product data sheet [16]. Therefore, it reinforces the reliability of the H1-1 emulator to approximate the corresponding device noise efficiently. There is a glitch that occurs at 76 hops, which results in a global increase in the measured negativity and fidelity afterwards. This is due to the long time gap between experiments performed before and after the 75-hop mark, where the noise parameters could have been optimised/totally different. From the plot, we also find that both dynamic circuit and post-selection approaches exhibit very similar behaviour regarding the amount of entanglement. This observation agrees with our expectations that the two approaches should only introduce little differences in the negativity and fidelity, given that their circuit depth differs by at most three gates. On the other hand, we have shown a distinct utility of the post-selection approach against dynamic circuits presented in previous work [6] using IBM quantum computers. This was a direct consequence of the accumulated latency for processing each feed-forward mid-circuit measurement sequentially and the relatively high level of noise in IBM quantum computers. From these results, we confirm that even after 56 hops, the negativity of the state teleported using a dynamic circuit can be preserved up to $0.291\pm 0.018$. When tested using the emulator, it still maintains a value of $0.224\pm 0.009$ up to 100 hops, which is just over 5 full cycles. Therefore, it is reasonable to expect that the actual limit of the number of hops the two-qubit state can teleport across using H1-1 quantum computer while preserving entanglement is far beyond 100.

### v3

@@ Abstract
Figure 3: Polar plots of the teleported two-qubit graph state showing (a) negativity and (b) fidelity on the radial axis versus the number of teleportation hops in the angular axis sampled from the H1-1 emulator averaged over four trials. The results for larger than 75 hops are marked using different colours to highlight the change in the measured entanglement due to different noise parameters arising from a time gap between experiments. The three data points marked in red represent results obtained from the physical H1-1 quantum computer for 9, 18 and 56 hops with respective negativities $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelities $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$, where experiments were conducted on 21 March, 25 March and 24 April 2024 respectively. Assuming the distribution is Gaussian, the half-width of the shaded strip is two times the standard error of the sampled negativity or fidelity to represent 95% confidence interval. On the other hand, the data obtained from real H1-1 quantum computers are only sampled in a single trial due to limited resources. Therefore, we use the technique of bootstrapping to resample the data from the measured sample distribution for 200 trials. The error bars represent 95% confidence interval of the bootstrapped negativities or fidelities. All results are mitigated with REM as indicated in Teleportation quality witnesses. Due to the feedforward gates applied based on mid-circuit measurements in the dynamic circuit approach, REM is applied to only the two qubits of the final teleported state. The corresponding plot with regular crossing axes are included in Figure 4.
@@
In Figure 3a and Figure 3b, we present the negativities and fidelities recorded from the real machine (red-marked data points) and the emulator (blue data), plotted in polar plots against the number of hops that the two-qubit graph states were teleported across using the teleportation strategies. The emulated data is averaged over four trials. Due to limited access to computational resources, experiments on the H1-1 quantum computer were conducted for only three scenarios of 9, 18, and 56 hops in a single trial of the QST experiment with 1000 shots for each QST circuit. We bootstrap 200 samples from each QST basis measurement for both negativities and fidelities. The error bars of the data obtained from the H1-1 quantum computer of size $\epsilon$ take the half-range of the 95% confidence interval of the bootstrapped negativities or fidelities, such that $\epsilon=\frac{1}{2}{\left(\left\lfloor a_{0.975N}\right\rfloor-\left\lceil a_{0.025N}\right\rceil\right)}$, where $\{a_{n}\},n=1,2,\cdots,N$ is the sequence of bootstrapped negativities or fidelities ordered from lowest to highest and $N=200$ is the sample size. For both the negativity and fidelity, we have used REM, the nearest physical probability vector, and the nearest physical density matrix algorithm as previously outlined to mitigate classical readout errors. After all, from the H1-1 quantum computer we obtain the negativity of $0.459\pm 0.009,0.388\pm 0.014,0.291\pm 0.018$ and fidelity of $0.958\pm 0.009,0.887\pm 0.014,0.791\pm 0.018$ corresponds to 9, 18 and 56 hops respectively. For the data obtained from the emulator, we report a clear trend that the entanglement of the teleported state decreases gradually, almost linearly, as the number of hops increases. This trend is closely in agreement with the data points obtained from the real device without significant deviations, as well as the trend of the two-qubit gate fidelity plot in the H1 model product data sheet [16]. Therefore, it reinforces the reliability of the H1-1 emulator to approximate the corresponding device noise efficiently. There is a clear change that occurs at 76 hops, which results in a global increase in the measured negativity and fidelity from this point. Given the long time gap between experiments performed before and after the 75-hop mark, this may indicate the changes in noise model parameters used in the emulator. From the plot, we also find that both dynamic circuit and post-selection approaches exhibit very similar behaviour regarding the amount of entanglement. This observation agrees with our expectations that the two approaches should only introduce little differences in the negativity and fidelity, given that their circuit depth differs by at most three gates. On the other hand, we have shown a distinct utility of the post-selection approach against dynamic circuits presented in previous work [6] using IBM quantum computers. This was a direct consequence of the accumulated latency for processing each feed-forward mid-circuit measurement sequentially and the relatively high level of noise in IBM quantum computers. From these results, we confirm that even after 56 hops, the negativity of the state teleported using a dynamic circuit can be preserved up to $0.291\pm 0.018$. When tested using the emulator, it still maintains a value of $0.224\pm 0.009$ up to 100 hops, which is just over 5 full cycles. Therefore, it is reasonable to expect that the actual limit of the number of hops the two-qubit state can teleport across using H1-1 quantum computer while preserving entanglement is far beyond 100.